Let T:P2 → p2 be defined by T(p(x)) = p(6x + 7) (a) Find the determinant of T. (b) Find the eigenvalues of T.
Answer #1

T (P(0.)=P(6x+7) let B={1, 2, na} Then 소 *) ri TW) = tex) = Then 6x +7 2 2 (6x+7) = 49+. 848x360 7 ET 0 6 49 84 36. B 0 0 The